# Coordinate-based Neural Network for Fourier Phase Retrieval

### v1

@@ Abstract
Fourier phase retrieval is essential for high-definition imaging of nanoscale structures across diverse fields, notably coherent diffraction imaging. This study presents the Single impliCit neurAl Network (SCAN), a tool built upon coordinate neural networks meticulously designed for enhanced phase retrieval performance. Bypassing the pitfalls of conventional iterative methods, which frequently face high computational loads and are prone to noise interference, SCAN adeptly connects object coordinates to their amplitude and phase within a unified network in an unsupervised manner. While many existing methods primarily use Fourier magnitude in their loss function, our approach incorporates both the predicted magnitude and phase, enhancing retrieval accuracy. Comprehensive tests validate SCAN’s superiority over traditional and other deep learning models regarding accuracy and noise robustness. We also demonstrate that SCAN excels in the ptychography setting.
@@ 1 Introduction
where $x$ is an $n$-dimensional spatial coordinate, and $k$ is a $n$-dimensional spatial frequency coordinate. $|F(k)|$ is what we measured on the detector while $\psi(k)$ is lost and needs to be recovered. The $f(x)$ and $|F(k)|$ in the above formulation could have different meanings in different applications. For example, in the transmission mode of CDI, the reconstructed amplitude and phase in $f(x)$ represent the imaginary and real parts of the sample’s refractive index; in the Bragg diffraction mode, its magnitude represents the electron density, and phase denotes the projection of the local deformations of the crystal lattice onto the reciprocal lattice vector Q of the Bragg peak about which the diffraction is measured [1].
@@
The intrinsic nature of objects, namely their coordinates, offers an accurate and compact representation. Each coordinate maps directly to its corresponding object amplitude and phase, forming the foundation of the Implicit Neural Representation (INR) concept, a paradigm that has recently been introduced to the research community [19, 20, 21]. Prior to our study, this framework demonstrated substantial success in image inversion tasks [22, 19, 23, 24, 25, 26]. For instance, NeRF [19] focuses on refining the object’s density and color based on multiple viewpoints from various angles. CoIL [24] trains a multilayer perceptron (MLP) to encode the complete measurement field and demonstrates superior performance in sparse-view computed tomography reconstruction. Notably, while DINER [27] and DNF [28] have applied successfully INR for phase recovery, a deeper dive reveals their observed measurements was in real space. To our knowledge, there remains an unexplored potential for the INR framework tailored for Fourier phase retrieval. Our contributions can be summarized as follows:
@@ 2 Methods
where $\tilde{M}\geq 2M-1,\tilde{N}\geq 2N-1$ provides additional information for image recovery. In ptychography, a localized probe is scanned over an extended object, and a series of measurements are collected from overlapping positions:
$I_{j}(\textbf{q})=|\mathcal{F}\{P*O(\textbf{r}_{j})\}|^{2},\vspace{-2pt}$ (3)
where ${}\mathcal{F}\{\cdot\}$ is the Fourier transform, $P$ denotes the probe, $O(\textbf{r})$ represents the object, $O(\textbf{r}_{j})$ is the part of the object scanned at position $j$, r is the real space, and q is the Fourier space. Our idea is to parameterize objects as a (continuous) function represented by $f_{\theta}(z)$, where $z$ is a set of coordinates, and $f_{\theta}(\cdot)$ symbolizes a deep neural network with parameters $\theta$. The following equations are formulated in the CDI context, with the ptychography variation conceptualized as scanning CDI employing $P$.
The network architecture is illustrated in Fig. 1. The first part is about image coordination, which involves encoding pixels using spatial coordinates. Subsequently, image coordinates are used as inputs for a fully connected network featuring sine activation [20] to obtain simultaneously the predicted phase and amplitude. The network parameters are iteratively adjusted during training using a meticulously defined loss function, see the details in Sec. 2.3.
@@ 2.3 Loss Function
2.3.1 Fourier Magnitude Loss ($\mathcal{L}_{m}$). Upon forward propagation through the network, the predicted amplitude and phase undergo additional forwarding to derive the predicted Fourier magnitude and phase. The primary loss metric evaluates the discrepancy between the predicted Fourier magnitude and the actual measurements. This form of loss assessment is commonly employed in unsupervised end-to-end methodologies [14, 15, 3, 16] because it directly quantifies the accuracy of the Fourier magnitude prediction. We write it as:
$\mathcal{L}_{m}=\left\||\mathcal{F}\{\widehat{O(\textbf{r})}\}|-\sqrt{I(\textbf{q})}\right\|_{2}^{2}.\vspace{-5pt}$ (5)
2.3.2 Distilled Fourier Phase Loss ($\mathcal{L}_{p}$). We further leverage the predicted phase by incorporating it into the loss calculation after suitable adjustments. Inspired by ER [7], which employs iterative processes between real and Fourier spaces, we integrate the predicted Fourier phase $u$ with measured Fourier magnitude $I$ to transition back to the real space, as illustrated in Fig. 1. Subsequently, we apply recognized constraints in the real space, denoted by $\odot S$, to this transformed object. These constraints enforce that the padding area should be zero and the object amplitude must reside within the 0-1 range. After imposing these constraints on the transformed object, we transition it forward to the Fourier space, deriving the Fourier magnitude for the second time. The secondary loss evaluates the disparity between this second-time predicted Fourier magnitude and the actual measurement.
The underlying rationale for the second loss $\mathcal{L}_{p}$ is that if the predicted Fourier phase is accurate, the merging of this phase with the true Fourier magnitude, coupled with the back-and-forth transitions between the Fourier and real spaces under established constraints, should yield a Fourier magnitude consistent with the measured one. The equations are detailed below. In this context, $\widehat{O_{a}}$ and $\widehat{O_{p}}$ represent the amplitude and the phase of the predicted object, respectively. They arise from the fusion of the measured magnitude $I$ and the predicted detector phase $u$.
@@
$\widehat{O_{a}}=|\mathcal{F}^{-1}\{Ie^{iu}\}|\odot S,\ \widehat{O_{p}}=angle(\mathcal{F}^{-1}\{Ie^{iu}\}),$ (7)
@@ 3.1 Experimental setup
Coherent Diffraction Imaging (CDI). In the existing studies, the oversample rate for most experiments is typically set around 2, with the actual object positioned ambiguously within the designated area. Consequently, a shifted symmetry may be observed during the recovery phase, as noted in [4]. Moreover, areas outside the object could potentially register values due to the quality of recovery. This scenario complicates the adoption of quantitative quality metrics like PSNR or SSIM. While the shrinkwrap method [29], which reduces minor values to zero, might provide a solution, the threshold for shrinkwrap can significantly vary across experiments, rendering comparisons unfair. Hence, we set the known object size to match the unpadded area in subsequent experiments to circumvent the shift issue and concentrate solely on the model’s recovery capability. We chose an oversample rate of approximately 5 for all benchmarked methods. Our simulated 2D object amplitude is a segment of the cameraman image ($50\times 50$ pixels), and the object phase derives from a portion of the brick image with the same size (Fig. 2).
@@ 3.2 Model Training
Table 3: Performance of phase retrieval methods under noise condition
@@ 4.1 Results of CDI setting
Noise Free Condition. Under noise-free conditions, Tab. 1 and Fig. 2 showed the best results for each method out of 3 random realizations (seeds). m-SiSPRNet achieved the best performance in amplitude recovery, followed by SCAN and wavelet. For object phase recovery, SCAN performed the best, then MLP and m-SiSPRNet. Phase is found to be harder than amplitude to recover. The same model could have quite different recovery abilities on these two properties. For instance, m-SiSPRNet’s phase recovery is 30 dB lower than its amplitude result. However, SCAN balanced the results and achieved good performance on both.
Stability. Tab. 2 showed the mean and standard deviation of 3 seeds’ experiment results under noise-free conditions. SCAN has the best amplitude and phase recovery performance when considering all seeds. It is noticed that decoder-encoder-based models, m-SiSPRNet and Double DIP, have a high variance in results. SCAN had a relatively lower variance, demonstrating its high stability.
@@
Table 4: Ablation study of different loss function.
@@
Time(s) 448.02 / / 324.52
Table 5: Ablation study of different model structures.
@@
ePIE 10.31/0.21 11.04/0.25 161.18 13.12/0.35 11.49/0.35 315.44 35.26/0.98 12.15/0.56 724.08
SCAN 25.17/0.82 19.40/0.77 45.21 27.94/0.93 24.59/0.89 72.24 36.69/0.998 35.65/0.98 341.76
Table 6: Ptychography reconstruction of ePIE and SCAN when the overlap rate is 30%, 50%, and 70%.
Ablation Study of Different Model Structures. Tab 5 and Fig 3 present several potential model implementations. The terms “clamp” and “shift” are utilized to denote alternative approaches to the Tanh activation function and absolute operation, respectively. Specifically, “clamp” refers to constraining the value within the range of $[-1,1]$, while “shift” involves scaling the value $x$ to the range of $[0,1]$ using the transformation $(x+1)/2$. Additionally, we explored utilizing two separate networks, one for object amplitude and the other for phase representation. The results demonstrate that our current implementation in SCAN outperforms all other approaches.

### v2

@@ Abstract
Fourier phase retrieval is essential for high-definition imaging of nanoscale structures across diverse fields, notably coherent diffraction imaging. This study presents the Single impliCit neurAl Network (SCAN), a tool built upon coordinate neural networks meticulously designed for enhanced phase retrieval performance. Remedying the drawbacks of conventional iterative methods which are easiliy trapped into local minimum solutions and sensitive to noise, SCAN adeptly connects object coordinates to their amplitude and phase within a unified network in an unsupervised manner. While many existing methods primarily use Fourier magnitude in their loss function, our approach incorporates both the predicted magnitude and phase, enhancing retrieval accuracy. Comprehensive tests validate SCAN’s superiority over traditional and other deep learning models regarding accuracy and noise robustness. We also demonstrate that SCAN excels in the ptychography setting.
@@ 1 Introduction
where $x$ is an $n$-dimensional spatial coordinate, and $k$ is an $n$-dimensional spatial frequency coordinate. $|F(k)|$ is what we measured on the detector while $\psi(k)$ is lost and needs to be recovered. The $f(x)$ and $|F(k)|$ in the above formulation could have different meanings in different applications. For example, in the transmission mode of CDI, the reconstructed amplitude and phase in $f(x)$ represent the imaginary and real parts of the sample’s refractive index; in the Bragg diffraction mode, its magnitude represents the electron density, and phase denotes the projection of the local deformations of the crystal lattice onto the reciprocal lattice vector Q of the Bragg peak about which the diffraction is measured [1].
@@
The intrinsic nature of objects, namely their coordinates, offers an accurate and compact representation. Each coordinate maps directly to its corresponding object amplitude and phase, forming the foundation of the Implicit Neural Representation (INR) concept, a paradigm that has recently been introduced to the research community [19, 20, 21]. Prior to our study, this framework demonstrated substantial success in image inversion tasks [19, 22, 23, 24, 25, 26]. For instance, NeRF [19] focuses on refining the object’s density and color based on multiple viewpoints from various angles. CoIL [24] trains a multilayer perceptron (MLP) to encode the complete measurement field and demonstrates superior performance in sparse-view computed tomography reconstruction. Notably, while DINER [27] and DNF [28] have successfully applied INR for phase recovery, a deeper dive reveals their objective is not Fourier phase retrieval. To our knowledge, there remains an unexplored potential for the INR framework tailored for Fourier phase retrieval. Our contributions can be summarized as follows:
@@ 2 Methods
where $\tilde{M}\geq 2M-1,\tilde{N}\geq 2N-1$ provide additional information for image recovery. In ptychography, a localized probe is scanned over an extended object, and a series of measurements are collected from overlapping positions:
$I_{j}(\textbf{q})=|\mathcal{F}\{P(\textbf{r}-\textbf{r}_{j})*O(\textbf{r})\}|^{2},\vspace{-2pt}$ (3)
where ${}\mathcal{F}\{\cdot\}$ is the Fourier transform, $O(\textbf{r})$ represents the object, $P(\textbf{r}-\textbf{r}_{j})$ denotes the probe scanning the object at position $j$, r and q are the real space and the Fourier space coordinate. Our idea is to parameterize objects as a (continuous) function represented by $f_{\theta}(z)$, where $z$ is a set of coordinates, and $f_{\theta}(\cdot)$ symbolizes a deep neural network with parameters $\theta$. The following equations are formulated in the CDI context, with the ptychography variation conceptualized as scanning CDI employing $P(\textbf{r}-\textbf{r}_{j})$.
The network architecture is illustrated in Fig. 1. The first part is about image coordination, which involves encoding pixels using spatial coordinates. Subsequently, image coordinates are used as inputs for a fully connected network featuring sine activation [20] to obtain simultaneously the predicted phase and amplitude. The network parameters are iteratively adjusted during training using a meticulously defined loss function; see the details in Sec. 2.3.
@@ 2.3 Loss Function
2.3.1 Fourier Magnitude Loss ($\mathbf{\mathcal{L}_{m}}$). Upon forward propagation through the network, the predicted amplitude and phase undergo additional forwarding to derive the predicted Fourier magnitude and phase. The primary loss metric evaluates the discrepancy between the predicted Fourier magnitude and the actual measurements. This form of loss assessment is commonly employed in unsupervised end-to-end methodologies [3, 14, 15, 16] because it directly quantifies the accuracy of the Fourier magnitude prediction. We write it as:
$\mathcal{L}_{m}=\left\||\mathcal{F}\{\widehat{O(\textbf{r})}\}|-\sqrt{I(\textbf{q})}\right\|_{2}^{2}.$ (5)
2.3.2 Distilled Fourier Phase Loss ($\mathbf{\mathcal{L}_{p}}$). We further leverage the predicted phase by incorporating it into the loss calculation after suitable adjustments. Inspired by ER [7], which employs iterative processes between real and Fourier spaces, we integrate the predicted Fourier phase $u$ with measured detector intensity $I$ to transition back to the real space, as illustrated in Fig. 1. Subsequently, we apply recognized constraints in the real space, denoted by $\odot S$, to this transformed object. These constraints enforce that the padding area should be zero and the object amplitude must reside within the 0-1 range. After imposing these constraints on the transformed object, we transition it forward to the Fourier space, deriving the Fourier magnitude for the second time. The secondary loss evaluates the disparity between this second-time predicted Fourier magnitude and the actual measurement.
The underlying rationale for the second loss $\mathcal{L}_{p}$ is that if the predicted Fourier phase is accurate, the merging of this phase with the true Fourier magnitude, coupled with the back-and-forth transitions between the Fourier and real spaces under established constraints, should yield a Fourier magnitude consistent with the measured one. The equations are detailed below. In this context, $\widehat{O_{a}}$ and $\widehat{O_{p}}$ represent the amplitude and the phase of the predicted object, respectively. They arise from the fusion of the measured detector intensity $I$ and the predicted detector phase $u$.
@@
$\widehat{O_{a}}=|\mathcal{F}^{-1}\{\sqrt{I}e^{iu}\}|\odot S,\ \widehat{O_{p}}=angle(\mathcal{F}^{-1}\{\sqrt{I}e^{iu}\}),$ (7)
@@ 3.1 Experimental setup
Coherent Diffraction Imaging (CDI). In the existing studies, the oversample rate for most experiments is typically set around 2, with the actual object positioned ambiguously within the designated area. Consequently, a shifted symmetry may be observed during the recovery phase, as noted in [4]. Moreover, areas outside the object could potentially register values due to the quality of recovery. This scenario complicates the adoption of quantitative quality metrics like PSNR or SSIM. While the shrinkwrap method [29], which reduces minor values to zero, might provide a solution, the threshold for shrinkwrap can significantly vary across experiments, rendering comparisons unfair. Hence, we set the known object size to match the unpadded area in subsequent experiments to circumvent the shift issue and concentrate solely on the model’s recovery capability. We chose an oversample rate of approximately 5 for all benchmarked methods. Our simulated 2D object amplitude is a segment of the cameraman image ($50\times 50$ pixels), and the object phase derives from a portion of the brick image of the same size (Fig. 2).
@@ 3.2 Model Training
Table 3: Performance of phase retrieval methods under noisy conditions (PSNR / SSIM)
@@ 4.1 Results of CDI setting
Noise Free Condition. Under noise-free conditions, Tab. 1 and Fig. 2 show the best results for each method out of 3 random realizations (seeds). m-SiSPRNet achieved the best performance in amplitude recovery, followed by SCAN and wavelet. For object phase recovery, SCAN performed the best, then MLP and m-SiSPRNet. Phase is found to be harder than amplitude to recover. The same model could have quite different recovery abilities on these two properties. For instance, m-SiSPRNet’s phase recovery is 30 dB lower than its amplitude result. However, SCAN balanced the results and achieved good performance on both.
Stability. Tab. 2 shows the mean and standard deviation of 3 seeds’ experiment results under the noise-free condition. SCAN has the best amplitude and phase recovery performance when considering all seeds. It is noticed that decoder-encoder-based models, m-SiSPRNet and Double DIP, have a high variance in results. SCAN had a relatively lower variance, demonstrating its high stability.
@@
Table 4: Ablation study of different loss function. (PSNR / SSIM)
@@
Time(s) 448.02 - - 324.52
Table 5: Ablation study of different model structures. (PSNR / SSIM)
@@
ePIE 9.75 / 0.20 11.03 / 0.20 76.58 10.08 / 0.17 11.14 / 0.23 141.10 29.70 / 0.90 17.72 / 0.80 394.69
SCAN 23.06 / 0.82 15.38 / 0.70 24.14 25.15 / 0.81 16.02 / 0.71 35.77 34.65 / 0.98 21.23 / 0.87 237.08
Table 6: Ptychography reconstruction of ePIE and SCAN when the overlap rate is 30%, 50%, and 70%. (PSNR / SSIM)
Ablation Study of Different Model Structures. Tab. 5 and Fig. 3 present several potential model implementations. The terms “clamp” and “shift” are utilized to denote alternative approaches to the Tanh activation function and absolute operation, respectively. Specifically, “clamp” refers to constraining the value within the range of $[-1,1]$, while “shift” involves scaling the value $x$ to the range of $[0,1]$ using the transformation $(x+1)/2$. Additionally, we explored utilizing two separate networks, one for object amplitude and the other for phase representation. The results demonstrate that our current implementation in SCAN outperforms all other approaches.